The sum of the distance of a point P from two perpendicular lines in a plane is 1, then the locus of P

Grade:12th pass

1 Answers

Y RAJYALAKSHMI
9 years ago
Let the perpendiculars lines be the x – axis and y – axis.  The sum of the distances from the the point P (x, y) is 1. ie.
|x| + |y| = 1  – this is the locus of the point P which would be the rhombus whose sides are 
x + y = 1; –x + y = 1; x – y = 1; –x – y = 1
If the perpendicular lines are other than co-ordinate axes, then we can get the same solution by using the transformation of the axes concept.